OpenStudy (issy14):

How do you calculate the Ac in a uniform circular motion in which the radius is infinitely large? V= 35m/s.

OpenStudy (theeric):

I would think this is very theoretical but we can look at from what we know. Centripetal acceleration is \(\dfrac {v^2} r\) If \(r\rightarrow \infty\), then \(a\rightarrow\dfrac{v^2}\infty\rightarrow 0\) It makes sense. The thing is hardly moving in a circle.

OpenStudy (anonymous):

@theEric is correct. Another way to look at this problem is to consider what happens to the radius when it become infinitely large. As the radius increases, its curvature decreases, and when the radius approaches infinity, the curvature approaches zero. So, at infinity, a curve with infinite radius becomes a straight line. There is no centripetal acceleration with movement in a straight line.

OpenStudy (issy14):

Thank you very much both of you, I wished my professor had explained that instead of just giving the answer. @PsiSquared & @theEric

OpenStudy (theeric):

You're welcome! It's too bad when teachers leave gaps in there explanations.. Sometimes it's better to repeat information than leave it out, in my opinion!! But OpenStudy can help with the gaps, I hope :) Good luck with your other studies!

OpenStudy (issy14):

@theEric yes it can, it has helped me a lot and it's good to interact with people that hopefully have the same passion as you do for the subject. Thanks again! good luck to you too =)

OpenStudy (theeric):

Cool! Thanks! :) Take care.